A company has a 11% WACC and is considering two mutually exclusive investments (that cannot be repeated) with the following cash flows:

0 1 2 3 4 5 6 7
Project A -$300 -$387 -$193 -$100 $600 $600 $850 -$180
Project B -$405 $135 $135 $135 $135 $135 $135 $0

The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the questions below.

Open spreadsheet

  1. What is each project's NPV? Round your answer to the nearest cent. Do not round your intermediate calculations.

    Project A: $  

    Project B: $  

  2. What is each project's IRR? Round your answer to two decimal places.

    Project A: %

    Project B: %

  3. What is each project's MIRR? (Hint: Consider Period 7 as the end of Project B's life.) Round your answer to two decimal places. Do not round your intermediate calculations.

    Project A: %

    Project B: %

  4. From your answers to parts a-c, which project would be selected?

    _________Project A or Project B

    If the WACC was 18%, which project would be selected?

    _________Project A or project B

  5. What is the cross over rate?

  6. Construct NPV profiles for Projects A and B. Round your answers to the nearest cent. Do not round your intermediate calculations. Negative value should be indicated by a minus sign.

    Discount Rate NPV Project A NPV Project B
    0% $   $  
    5 $   $  
    10 $   $  
    12 $   $  
    15 $   $  
    18.1 $   $  
    24.29 $   $  
  7. Calculate the crossover rate where the two projects' NPVs are equal. Round your answer to two decimal places. Do not round your intermediate calculations.

    %

  8. What is each project's MIRR at a WACC of 18%? Round your answer to two decimal places. Do not round your intermediate calculations.

    Project A: %

    Project B: %
Answer #1

a.

Time Years 0 1 2 3 4 5 6 7
Project A -300 -387 -193 -100 600 600 850 -180
Project B -400 133 133 133 133 133 133
Discounting Rate 1.11 1 0.900901 0.811622 0.731191 0.658731 0.593451 0.534641 0.481658
Discounted Flows A -300 -348.649 -156.643 -73.1191 395.2386 356.0708 454.4447 -86.6985
Discounted Flows B -400 119.8198 107.9458 97.24845 87.61122 78.92903 71.10723 0
NPV Project A 240.6447
Project B 162.6615

b. Internal Rate of Return is the rate at which cash outflows is equal to cash inflows. The different rates are tried and then, we find the IRR for Project B is 24.18% and Project A is 18.10%

Time Years 0 1 2 3 4 5 6 7
Project B -400 133 133 133 133 133 133
Discounting Rate 1.2418 1 0.805283 0.64848 0.52221 0.420527 0.338643 0.272703 0.219603
Discounted Flows B -400 107.1026 86.24786 69.45391 55.93003 45.03948 36.26951 0
NPV Project B 0.043376
Time Years 0 1 2 3 4 5 6 7
Project A -300 -387 -193 -100 600 600 850 -180
Discounting Rate 1.181 1 0.84674 0.716969 0.607086 0.514044 0.435262 0.368554 0.312069
Discounted Flows A -300 -327.688 -138.375 -60.7086 308.4265 261.1571 313.2705 -56.1724
NPV Project A -0.09034

c. Modified Internal rate of Return is future value of positive cash flows divided by the present value of cash outflows. Hence MIRR of A = (600+800+850)/(300+328+138+61+57)*100 = 2.32%

MIRR of B= (133+133+133+133+133+133+0)/400*100 = 1.99%

d. From Answers (a-c) ,we would accept Project A since it has a better NPV and also a higher MIRR than Project B.

If the WACC was 18%, then the NPV of Project A is $2 and the NPV of Project B is $66. Hence, we would accept Project B in that case.

The excel calculations are also added.Time Years Project A Project B Discounting Rate Discounted Flows A Discounted Flows B -300 -400 1 -300 -400 1 3 4 56 -387 -19